Annapolis Company has two service departments (Computer Operations & Maintenance Services). Annapolis has two production departments (Mixing Department & Packaging Department.) Annapolis uses a step allocation method where the Computer Operations Department is allocated to all departments and Maintenance Services is allocated to the production departments.  All allocations are based on total employees. Computer Operations has costs of $100,000 and Maintenance Services has costs of $120,000 before any allocations. What amount of Maintenance Services total cost is allocated to the Mixing Department? (round to closest whole dollar) Employees are:

            Computer Operations 2

            Maintenance Services 2

Mixing Department 7

Packaging Department 7
Answer #1

The amount of maintenance services total cost allocated to mixing department :

Computer operations  cost allocated to maintenance services = $100,000 * 2/16 = $12,500

Total cost of maintenance services to be allocated = $120,000 + 12,500 = $132,500

Amount of maintenance services total cost allocated to mixing department = $132,500 * 7/14 = $66,250